caluculate the Resolution of electron microscopes and optical microscopes


1
Expert's answer
2021-02-22T10:25:18-0500

For electron microscopes:

The de Broglie wavelength associated with a particle having momentum p is given as follows.

"λ = \frac{h}{p}"

Substitute "\frac{\sqrt{K^2+2KE_0}}{c}" for p.

"λ = \frac{h}{\frac{\sqrt{K^2+2KE_0}}{c}} \\

= \frac{hc}{\sqrt{K^2+2KE_0}}"

For a 3.0-MV transmission electron microscope:

"K = 3.0 \times 10^6 \;eV \\

E_0 = 0.511 \times 10^6 \;eV \\

λ = \frac{1240 \;eV \times nm}{\sqrt{(3.0 \times 10^6 \;eV)^2+2(3.0 \times 10^6 \;eV)(0.511 \times 10^6 \;eV)}} \\

= 3.57 \times 10^{-12} \;m \\

= 3.57 \;pm"

For optical microscopes:

"d= \frac{λ}{2NA}"

d is the resolution

λ is the wavelength

NA is the numerical aperture, which gathers light and resolve the specimen.

Where λ is the wavelength of light used to image a specimen. If using a green light of 514 nm and an oil immersion objective with an NA of 1.45, then the (theoretical) limit of resolution will be 177 nm.

"d= \frac{514}{2 \times 1.45} = 177 \;nm"